OpenStudy (anonymous):

Find an equation of the line having the given slope and containing the given point. m=5/6 (8,-4) The equation of the line is y= ?

zepdrix (zepdrix):

Oh they want it in terms of y=? Ok I guess we should use slope-intercept form then.

zepdrix (zepdrix):

Slope-Intercept Form of a Line:\[\Large\rm y=\color{orangered}{m}x+b\] And they gave us this information:\[\Large\rm \color{orangered}{m=\frac{5}{6}}, \qquad\qquad (x,y)=(8,~-4)\]

zepdrix (zepdrix):

We can plug our slope value directly in,\[\Large\rm y=\color{orangered}{\frac{5}{6}}x+b\]But we still need a b value ( the y-intercept ). To find that, we plug in our coordinate pair and try to isolate the b term.

zepdrix (zepdrix):

Plugging in the coordinate pair gives us,\[\Large\rm -4=\color{orangered}{\frac{5}{6}}\cdot 8+b\]Understand how to solve for the b?

OpenStudy (anonymous):

cant see any posting any pictures or drawing at all

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

You would use point slope- form

OpenStudy (anonymous):

well i put y=5/6x-4

OpenStudy (anonymous):

I got y = 5/6x-32/3... I'm not sure if it's correct

OpenStudy (mathmale):

Were I doing this problem, I'd definitely use the point-slope form of the equation of a straight line first. We know that m=(5/6) and one point is (8,-4); this is given. Substitute this information into the point-slope form of the equation of a straight line. Let me know if you need further help with this.

OpenStudy (anonymous):

try slope-intercept so y = mx + b y being b and m being slope y = 5/6x - 4